# Uncertainty-Aware Hardware Trojan Detection Using Multimodal Deep Learning

### 15 Jan 2024 v1

@@ Abstract
The risk of hardware Trojans being inserted at various stages of chip production has increased in a zero-trust fabless era. To counter this, various machine learning solutions have been developed for the detection of hardware Trojans. While most of the focus has been on either a statistical or deep learning approach, the limited number of Trojan-infected benchmarks affects the detection accuracy and restricts the possibility of detecting zero-day Trojans. To close the gap, we first employ generative adversarial networks to amplify our data in two alternative representation modalities, a graph and a tabular, ensuring that the dataset is distributed in a representative manner. Further, we propose a multimodal deep learning approach to detect hardware Trojans and evaluate the results from both early fusion and late fusion strategies. We also estimate the uncertainty quantification metrics of each prediction for risk-aware decision-making. The outcomes not only confirms the efficacy of our proposed hardware Trojan detection method but also opens a new door for future studies employing multimodality and uncertainty quantification to address other hardware security challenges.
@@
Hardware Trojan, Multimodal Deep Learning, Uncertainty Quantification
@@ I Introduction
Hardware Trojans (HT) have been a growing concern in today’s fabless semiconductor manufacturing, as malicious modifications can be made by attackers for a variety of reasons, such as information leakage, incorrect operation, or inflicting damage on the chip [1, 2, 3, 4]. While comprehensive approaches are vital for countering HTs, they entail certain drawbacks. Formal methods and simulation-based testing can be resource-intensive and time-consuming. Intrusion detection systems may yield false alarms, disrupting operations. Establishing a secure supply chain can limit flexibility in supplier selection.
Recently, Machine Learning (ML) has emerged as a powerful tool for detecting HTs [5, 6, 7, 8, 9]. It leverages algorithms to identify intricate patterns indicative of Trojans, even in increasingly sophisticated attacks. By training on diverse datasets, ML models can classify circuits as Trojan-free or Trojan-infected. Real-time processing enables continuous monitoring and immediate threat response. However, challenges exist, for example, acquiring large and diverse datasets, especially for rare Trojans, which can be difficult. Moreover, models are susceptible to adversarial attacks [10], potentially undermining their decision-making. Interpretability [11] and explainability [12] are crucial for trust but can be complex in this context. Additionally, resource-intensive training and deployment may limit accessibility for smaller manufacturers. Continuous retraining is necessary to adapt to evolving Trojan techniques, adding complexity to maintenance.
@@ I-A Related Works
Multimodal Deep Learning (DL) has been a well-explored topic in the Artificial Intelligence (AI) community. Early research, exemplified by Deep Boltzmann Machines (DBM) focused on the model’s capacity to understand probability distributions across inputs with multiple modes [18]. Additionally, applications of uncertainty-aware multimodal learning [19] have been successfully demonstrated in healthcare [20] and in scenarios involving multimodal task distributions [21], particularly in safety-critical environments. In our study, we specifically target the fusion of graph [22, 23] and Euclidean data as the modalities of interest along with uncertainty estimation.
Moreover, when working in the hardware security domain, it is expected to have fewer data points that are malicious or Trojan-infected. In this context, it becomes a necessity to work with small data [24] and the same has been achieved in various domains such as material science [25], and also for anomaly detection [26].
@@ I-B Contributions
In this paper, we investigate the feasibility of applying a multimodal ML approach for HT identification by deriving two data representations of circuits. The first is graphical representation [27] of circuits, and the second is euclidean data [28] derived by processing the Abstract Syntax Tree (AST) of the RTL files (Verilog). Although the use of multimodal approaches for improved model accuracy has been used in other domains, we do not see any application in Trojan identification. For uncertainty-aware multimodal learning, we believe the logic should be implemented at the information fusion level of the modalities, and for this, we leverage the $p$-values aggregation with conformal prediction. Our main contributions are as follows:
@@
  • Suggesting a model fusion approach using $p$-values with uncertainty quantifier. By employing $p$-values as a statistical measure, we can systematically assess the significance of each modality’s contribution to the overall prediction. This not only enhances the interpretability of the fusion process but also enables more robust decision-making.
  • We address the challenges of missing modalities and solve the issue of handling imbalanced and small dataset by leveraging generative adversarial networks.
@@ II-A Multimodal Learning
Multimodal learning [29] addresses complex problems by integrating information from multiple modalities, such as text, images, and audio, to obtain a comprehensive understanding of a given phenomenon. In our case, we used graphical data and tabular representations of the source circuits. This approach enables models to capture nuanced relationships that may be overlooked when considering each modality in isolation. The fusion of information from different sources empowers the model to make more accurate and robust predictions.
@@ III Multimodal Hardware Trojan Detection
While state-of-the-art works on HT detection have focused mainly on choosing the right algorithm and choosing different representations of the dataset for improved accuracy, incorporating different modalities of the same data and feeding it to the ML system has not been investigated. By performing information fusion derived from different modalities, a more refined data representation can be achieved. Furthermore, in a real-time scenario, we encounter missing values while collecting data, and this may lead to missing modalities when dealing with a multimodal ML approach. So, we also need a method that handles missing modalities for any given dataset. Lastly, in the domain of hardware security, it is difficult to get enough data for training, especially the labels marked as Trojan-infected because of the rarity of the event. In such a situation, we need to work with limited data.
@@
After obtaining a sufficient number of data points for the experiment, we implement multimodal ML using the graph and tabular data. Specifically, we have employed a CNN for binary classification. It’s worth mentioning that any ML model can be optimized through hyper-parameter tuning to enhance accuracy. However, our primary emphasis is on assessing the effectiveness of uncertainty-aware multimodality by accessing early and late fusions. Finally, the model will be used to make further informed decisions for the detection of HTs.
@@ V Conclusion
In this study, we have addressed the growing concern of maliciously inserted hardware Trojans into chips at various stages of production in an era where fabless manufacturing is hard to trust. To tackle this issue, we adopted an innovative approach by utilizing generative adversarial networks to expand our dataset with two distinct representation modalities: graph and tabular. Additionally, we introduced an uncertainty-aware multimodal deep learning framework called NOODLE for detecting hardware Trojans. We assessed our findings using both early and late fusion strategies, offering a comprehensive evaluation of our approach’s efficacy. Moreover, we integrated metrics for uncertainty quantification for each prediction, enabling us to make decisions that are mindful of potential risks. The utilization of multimodality and uncertainty quantification shows great potential for addressing other critical challenges in hardware security. These contributions collectively represent a significant step forward in enhancing the security and reliability of hardware systems in the face of emerging threats.
@@ Acknowledgment
This work is supported by the National Science Foundation under Award No. 2245247.

### 23 Jan 2024 v2

@@ Abstract
The risk of hardware Trojans being inserted at various stages of chip production has increased in a zero-trust fabless era. To counter this, various machine learning solutions have been developed for the detection of hardware Trojans. While most of the focus has been on either a statistical or deep learning approach, the limited number of Trojan-infected benchmarks affects the detection accuracy and restricts the possibility of detecting zero-day Trojans. To close the gap, we first employ generative adversarial networks to amplify our data in two alternative representation modalities: a graph and a tabular, which ensure a representative distribution of the dataset. Further, we propose a multimodal deep learning approach to detect hardware Trojans and evaluate the results from both early fusion and late fusion strategies. We also estimate the uncertainty quantification metrics of each prediction for risk-aware decision-making. The results not only validate the effectiveness of our suggested hardware Trojan detection technique but also pave the way for future studies utilizing multimodality and uncertainty quantification to tackle other hardware security problems.
@@
Hardware Trojan; Multimodal Deep Learning; Uncertainty Quantification
@@ I Introduction
Hardware Trojan (HT) insertion has become a major concern in today’s fabless semiconductor manufacturing since attackers can make malicious modifications for a variety of reasons, such as information leakage, incorrect operation, or inflicting damage on the chip [1, 2, 3, 4]. While comprehensive approaches are vital for countering HTs, they entail certain drawbacks. Formal methods and simulation-based testing can be resource-intensive and time-consuming. Intrusion detection systems may yield false alarms, disrupting operations. Establishing a secure supply chain can limit flexibility in supplier selection.
Recently, Machine Learning (ML) has emerged as a powerful tool for detecting HTs [5, 6, 7, 8, 9]. It leverages algorithms to identify intricate patterns indicative of Trojans, even in increasingly sophisticated attacks. By training on diverse datasets, ML models can classify circuits as Trojan-free or Trojan-infected. Real-time processing enables continuous monitoring and immediate threat response. However, challenges exist, for example, acquiring large and diverse datasets, especially for rare Trojans, which can be difficult. Moreover, models are susceptible to adversarial attacks [10], potentially undermining their decision-making. Interpretability [11] and explainability [12] are crucial for trust but can be complex in this context. Additionally, resource-intensive training and deployment may limit accessibility for smaller manufacturers. Continuous retraining is necessary to adapt to evolving Trojan techniques [13], adding complexity to maintenance.
@@ I-A Related Works
Multimodal Deep Learning (DL) has been a well-explored topic in the Artificial Intelligence (AI) community. Early research, exemplified by Deep Boltzmann Machines (DBM) focused on the model’s capacity to understand probability distributions across inputs with multiple modes [19]. Additionally, applications of uncertainty-aware multimodal learning [20] have been successfully demonstrated in healthcare [21] and in scenarios involving multimodal task distributions [22], particularly in safety-critical environments. In our work, we target the fusion of graph [23, 24] and Euclidean data as the modalities of interest along with uncertainty estimation.
Moreover, when working in the hardware security domain, it is expected to have fewer data points that are malicious or Trojan-infected. In this context, it is necessary to work with small data [25] and this has been achieved in various domains such as material science [26] and anomaly detection [27].
@@ I-B Contributions
In this paper, we investigate the feasibility of applying a multimodal ML approach for HT identification by deriving two data representations of circuits. The first is graphical representation [28] of circuits, and the second is euclidean data [29] derived by processing the Abstract Syntax Tree (AST) of the RTL files (Verilog). Although the use of multimodal approaches for improved model accuracy has been used in other domains, we do not see any application in Trojan identification. For uncertainty-aware multimodal learning, we believe the logic should be implemented at the information fusion level of the modalities, and thus, we leverage the $p$-values aggregation with conformal prediction. Our main contributions are as follows:
@@
  • Suggesting a model fusion approach using $p$-values with an uncertainty quantifier. By employing $p$-values as statistical measures, we can systematically assess each modality’s contribution to the overall prediction. This not only enhances the interpretability of the fusion process but also enables more robust decision-making.
  • Addressing the challenges of missing modalities and solving the issue of handling an imbalanced and small dataset by leveraging generative adversarial networks.
@@ II-A Multimodal Learning
Multimodal learning [30] addresses complex problems by integrating information from multiple modalities, such as text, images, and audio, to obtain a comprehensive understanding of a given phenomenon. In our case, we use graphical data and tabular representations of the source circuits. This approach enables models to capture nuanced relationships that may be overlooked when considering each modality in isolation, and thus empowers the model to make more robust predictions.
@@ III Multimodal Hardware Trojan Detection
While state-of-the-art works on HT detection have focused mainly on choosing the right algorithm and choosing different representations of the dataset for improved accuracy, incorporating different modalities of the same data and feeding it to the ML system has not been investigated. By performing information fusion derived from different modalities, a more refined data representation can be achieved. Furthermore, in a practical scenario, we encounter missing values while collecting data, and this may lead to missing modalities when dealing with a multimodal ML approach. So, we also need a method that handles missing modalities for any given dataset. Lastly, in the domain of hardware security, it is difficult to get enough data for training, especially the labels marked as Trojan-infected because of the rarity of the event. In such a situation, we need to work with limited data.
@@
After obtaining a sufficient number of data points for the experiment, we implement multimodal ML using the graph and tabular data. Specifically, we have employed a CNN for binary classification. It is worth mentioning that any ML model can be optimized through hyperparameter tuning to enhance accuracy. However, our primary emphasis is on assessing the effectiveness of uncertainty-aware multimodality by accessing early and late fusions. Finally, the model will be used to make further informed decisions for the detection of HTs.
@@ V Conclusion
In this paper, we have addressed the growing concern of maliciously inserted hardware Trojans into chips at various stages of production in an era where fabless manufacturing is hard to trust. Specifically, we adopted an innovative approach by utilizing generative adversarial networks to expand our dataset with two distinct representation modalities: graph and tabular. Additionally, we introduced an uncertainty-aware multimodal deep learning framework called NOODLE for detecting hardware Trojans. We assessed our findings using both early and late fusion strategies, offering a comprehensive evaluation of our approach’s efficacy. Moreover, we integrated metrics for uncertainty quantification for each prediction, enabling us to make decisions that are mindful of potential risks. The utilization of multimodality and uncertainty quantification shows great potential for addressing other critical challenges in hardware security such as logic locking [41, 42, 43, 44]. These contributions collectively represent a significant step forward in enhancing the security and reliability of hardware systems in the face of emerging threats.
@@ Acknowledgment
This material is based upon work supported by the National Science Foundation under Award No. 2245247.